Two 20.0-g ice cubes at -14.0 degree C are placed
Answer #1

NOTE: We consider rise or fall in temperature. So X oC rise or fall = X kelvin rise or fall

Two 20.0 g ice = 40g

Molar mass of water = 18 g/mol

So, 18 g of water = 1 mol

40 g of water = (40/22) mol = 2.22 mol

Heat required to heat ice from -14 oC to 0 oC

= Specific heat(H2O solid) * Mass * Temp rise

= (37.7 J mol-1 K-1) * (2.22 mol) * (14 K)

= 1171.72 J

Heat required to melt ice at 0 oC to water at 0 oC

= Mass * Enthalpy of fusion

= (2.22 mol) * (6.01 kJ mol-1)

= (2.22 mol) * (6010 J mol-1)

= 13342.2 J

Call final temperature X.

Heat required to heat water at 0 oC to X oC

= Mass * specific heat H2O (liquid) * Temp rise

= (2.22 mol) * (75.3 J mol-1 K-1) * (X K)

= 167.2X J

Total heat gained by ice = 1171.72 J + 13342.2 J + 167.2X

                                       = 14513.92 + 167.2X

Total heat lost by water = Mass * Specific heat * Temp Fall

                                       = (295/18 mol) * (75.3 J mol-1 K-1) * (25 - x)

                                       = (16.39 mol) * (75.3 J mol-1 K-1) * (25 - x)

                                       = 1234.2(25 - X)

                                       = 30855 - 1234.2X

Total heat gained by ice = Total heat lost by water

14513.92 + 167.2X = 30855 - 1234.2X

167.2X + 1234.2X = 30855 - 14513.92

1401.4X = 16341.08

X = 11.66